OpenStudy (anonymous):

(1+tan^2x)sin^2x=tan^2x

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle (1+\tan^2x)\sin^2x=\tan^2x }\) \(\large\color{black}{ \displaystyle (\sec^2x)\sin^2x=\tan^2x }\) \(\large\color{black}{ \displaystyle \frac{1}{\cos^2x}\sin^2x=\frac{\sin^2x}{\cos^2x} }\)

OpenStudy (solomonzelman):

and then you see it from there